Project

General

Profile

NDOS Commissioning Meeting 2011/08/25

Attendance: Jon, Rafael, Gavin, Enrique, Carl B., Zukai, Minerba, Gary, Susan, Nick, Chad, Martin, Leon

Agenda:

  1. Beam & Detector Update
  2. DataCheck Update
  3. DAQ Update
  4. DCS Update
  5. AOB

Beam & Detector Update

  • Next Tuesday, Aug. 30, power will be turned off at the NDSB at 9 am for about an hour for grounding studies.
  • Leon:
    • ~20 APDs have passed all tests at CalTech
    • Looking at heatsinks on APDs removed last week, there is clearly a sealing issue with the heatsinks. Pulled a vacuum on the seals, and the vacuum failed in seconds (compared to hours seen on "good" APDs).
    • Before the good APDs are sent back to FNAL, would like to address this sealing issue, in case the problem is on the APD carrier board (eg).

DataCheck Update

  • Plots from the past week look pretty good. There are some normalization issues with some of the plots that will be addressed soon.
  • "Good run" cuts investigated by Susan. One curious thing she's observed in many of the distributions is a dual-peak structure which seems to correlate with time; one peak corresponds to data collected after May, the other is from data collected before. She is investigating into the root cause.
  • The nearline is up and running; shifters can go to the nearline website to see plots from the last 24 hours. One concern is that data are being copied to the /data2/ disk only once per day; this needs to be done more frequently. Checklist plots are a work-in-progress.

DCS Update

  • Martin:
    • code for alarms for FEB temp. and drive current in IOC is ready, and IOCs will be updated tomorrow morning.
    • should the GUI be updated tomorrow too? That means that no DCS experts will be around over the weekend to deal with any issues. Jon: try it out tomorrow and see how it goes. If it looks to behave reasonably, then leave it for the weekend. Otherwise, wait until Monday.

DAQ Update

  • Peter:
    • Andrew N. was able to successfully take control of the NDOS DAQ and run it from the Ash River CR.
    • Ron fixed the problem with external timing commands on DCMs. Jon will remove instructions to shifters to "force" ET.

AOB